OpenStudy (anonymous):

can someone give me the difference of dot product to cross product of vector???

OpenStudy (zehanz):

The dot product is just a number. The cross product is another vector.

pooja195 (pooja195):

Dot Product in R3 If v = ai + bj + ck and w = di + ej + fk then v . w = ad + be + cf

OpenStudy (zehanz):

Cross product: (bf-ce)i + (cd-af)j + (ae-bd)k

OpenStudy (anonymous):

thanks for the answer. then can a cross product be a form of matrix and the answer could have been in a form of i + j + k...??? just like that hehehe

OpenStudy (amistre64):

cross product does have a certain determinant of a matrix quaility to it yes

OpenStudy (anonymous):

okey then, why does the dot product also called as the scalar product. does it always have a scalar.

OpenStudy (amistre64):

the dot product is like matrix multiplication; it doesnt form a new vector .... but a scalar

OpenStudy (zehanz):

Because it is just a number. If that is the case it is called a scalar (only has magnitude). If a variable has a magnitude AND a direction, it is called a vector.

OpenStudy (anonymous):

ohhhhhh.... k. tnx :] hehehe